# The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069

### v4

@@ 4.1 Unconservative mass transfer
It is also possible for material to be lost through the Lagrangian point L2, and this is more likely to happen for systems with high eccentricity (Sepinsky et al., 2007a). This lost material may form a circumbinary binary disk, which may significantly affect the orbital evolution(Dermine et al., 2013). I will not discuss this scenario here, but rather the case where this material leaves the binary system altogether.
Differentiating the total angular momentum gives
@@
Assuming the mass transfer occurs only instantaneously at the periastron, then $\dot{r}_{p}=0$. The fraction of mass that is accreted to the central black hole is defined as $\gamma=-\dot{M}_{h}/\dot{M}$. The mass lost by the system also carries away angular momentum. Defining that the ratio of its specific angular momentum to that of the stripped star is $\lambda$, the change of the total angular momentum is $\dot{J}/J=\lambda\dot{M}_{t}/qM_{t}$. Then Eq. (25) becomes
@@
When $\gamma=1$, they reduce to Eq. (5) and (7).
@@
It is about half of what it is at $\gamma=1$, as is $\langle\dot{a}/a\rangle_{\dot{M}}$. The change of $\beta$ now is
@@
In this scenario, the effect of mass transfer is about 2 times the effect of the gravitational radiation in the balance state by comparing it with the case of $\gamma=1$ as shown in Fig. 2. So the effect of mass transfer is still dominant on the orbital evolution in the balance state for the main-sequence star.
@@ 4.2 Extreme mass ratio inspirals
Extreme mass ratio inspirals are the potential gravitational wave sources, and they are expected to form when compact objects gradually inspiral via gravitational wave radiation after being captured in a high eccentricity orbit (Amaro-Seoane, 2018). Our results show that if the mass transfer is triggered during this process, the inspiral will be terminated and thus fail to become a gravitational wave source.
4.3 On the other QPEs
GSN 069 is not the only one galaxy found to have QPEs (Giustini et al., 2020; Song et al., 2020; Arcodia et al., 2021; Chakraborty et al., 2021), but is the only one with both TDE and QPEs (Shu et al., 2018; Sheng et al., 2021; Miniutti et al., 2023b). The orbiting WD is likely captured from a binary via the Hills mechanism (Hills, 1988; Wang et al., 2022), while the TDE possibly originates from a disruption of a common envelope (Wang, 2023). In this situation, the fallback disk is coplanar with the orbit of the common envelope, which is almost coplanar with the orbit of the captured WD. Thus, the orbiting star moves inside the disk for GSN 069.
As for the other QPEs sources, the situation may be different. The pre-existing disk may not be present, or the pre-existing disk may be inclined to the orbit of the stripped star. In the former case, the orbital evolution is the same as discussed in Section 2. In the latter case, the effect of the disk on the orbit will be different from that discussed in Section 3. For example, Linial & Quataert (2024) found the hydrodynamic drag force received by the star as it passes through the inclined disk can make the orbital period decrease.
The TDE in GSN 069 makes it unique. On the one hand, it may be the only source whose accretion disk is coplanar with the orbit of the stripped star. On the other hand, the accretion rate of its disk is dominated by the TDE and therefore evolves rapidly. These two points lead to a complicated orbital evolution, making the QPEs in GSN 069 an excellent source to study the interaction between the disk and the object moving inside it.
@@ 5 Summary
By accounting for the effect of the mass transfer, the evolutionary path of the stripped star is different from the evolutionary only with gravitational wave radiation. The criterion of the unstable mass transfer is changed and the fate of the stripped star is to escape the SMBH rather than be disrupted by it. In this case, however, mass transfer remains unstable.
The TDE in GSN 069 may be a disruption of the common envelope before the WD was captured, so the orbiting WD is likely moving inside the fallback disk. The hydrodynamic drag force by the disk can effectively reduce $\beta$ and stabilize the mass transfer, which can explain the observed long term evolution in the intensity of QPEs. After the TDE re-brightened, the mass transfer increases due to the weakening of the drag force caused by the reduced density of the inner zone of the accretion disk.As the accretion rate of the fallback disk gradually decreases, the drag force by the disk will decrease until it is negligible. Then the mass transfer will start to be unstable eventually the WD will escape the SMBH.
The spectrum evolution of QPEsis also directly related to the evolution of the accretion disk. Future work will require further quantitative analysis of the evolution of mass transfer in combination with specific radiative processes. And continued observations of GSN 096 will help us to understand the interaction of the disk with the star inside it and to test our model.

### v5

@@ 4.1 Unconservative mass transfer
It is also possible for material to be lost through the Lagrangian point L2 and this is more likely to happen for systems with high eccentricity (Sepinsky et al., 2007a). This lost material may form a circumbinary binary disk, which may significantly affect the orbital evolution(Dermine et al., 2013). This scenario is not discussed here, in favor of focusing on the case where this material leaves the binary system altogether.
Differentiating the total angular momentum gives:
@@
Assuming the mass transfer occurs only instantaneously at the periastron, then $\dot{r}_{p}=0$. The fraction of mass that is accreted to the central black hole is defined as $\gamma=-\dot{M}_{h}/\dot{M}$. The mass lost by the system also carries away angular momentum. Based on the assumption that the ratio of its specific angular momentum to that of the stripped star is $\lambda$, the change in the total angular momentum is $\dot{J}/J=\lambda\dot{M}_{t}/qM_{t}$. Then Eq. (25) becomes:
@@
When $\gamma=1$, these are reduced to Eq. (5) and (7).
@@
This result is about half of what it is at $\gamma=1$, as is $\langle\dot{a}/a\rangle_{\dot{M}}$. The change in $\beta$ now is:
@@
In this scenario, the effect of mass transfer is about twice the effect of the gravitational radiation in the balance state by comparing it with the case of $\gamma=1,$ as shown in Fig. 2. Thus, the effect of mass transfer is still dominant on the orbital evolution in the balance state for the main sequence star.
@@ 4.2 Extreme mass ratio inspirals
Extreme mass-ratio inspirals are potential gravitational wave sources and they are expected to form when compact objects gradually inspiral via gravitational wave radiation after being captured on a highly eccentricity orbit (Amaro-Seoane, 2018). Our results show that if the mass transfer is triggered during this process, the inspiral will be terminated and will thus fail to become a gravitational wave source.
4.3 Considering the other QPEs
GSN 069 is not the only galaxy found to have QPEs (Giustini et al., 2020; Song et al., 2020; Arcodia et al., 2021; Chakraborty et al., 2021), but is the only one with both TDEs and QPEs (Shu et al., 2018; Sheng et al., 2021; Miniutti et al., 2023b). The orbiting WD is likely captured from a binary via the Hills mechanism (Hills, 1988; Wang et al., 2022), while the TDE possibly originates from a disruption of a common envelope (Wang, 2024). In this situation, the fallback disk is coplanar with the orbit of the common envelope, which is almost coplanar with the orbit of the captured WD. Thus, the orbiting star moves inside the disk for GSN 069.
As for the other QPE sources, the situation may be different. The pre-existing disk may not be present or the pre-existing disk may be inclined to the orbit of the stripped star. In the former case, the orbital evolution is the same as discussed in Section 2. In the latter case, the effect of the disk on the orbit will be different from that discussed in Section 3. For example, Linial & Quataert (2024) found the hydrodynamic drag force received by the star as it passes through the inclined disk can make the orbital period decrease.
The TDE in GSN 069 makes this system unique. On the one hand, it may be the only source whose accretion disk is coplanar with the orbit of the stripped star. On the other hand, the accretion rate of its disk is dominated by the TDE and, therefore, it evolves rapidly. These two points lead to a complicated orbital evolution, making the QPEs in GSN 069 an excellent source for studying the interaction between the disk and the object moving inside it.
@@ 5 Summary
By accounting for the effect of the mass transfer, the evolutionary path of the stripped star is different from the evolutionary only with gravitational wave radiation. The criterion of the unstable mass transfer is changed and the fate of the stripped star is to escape the SMBH, rather than be disrupted by it. In this case, however, mass transfer remains unstable.
The TDE in GSN 069 may be a disruption of the common envelope before the WD was captured, so the orbiting WD is likely moving inside the fallback disk. The hydrodynamic drag force by the disk can effectively reduce $\beta$ and stabilize the mass transfer, which may explain the observed long-term evolution in the intensity of QPEs. After the TDE is re-brightened, the mass transfer increases due to the weakening of the drag force caused by the reduced density of the inner zone of the accretion disk. As the accretion rate of the fallback disk gradually decreases, the drag force by the disk will decrease until it is negligible. Then the mass transfer will start to be unstable the WD eventually escapes the SMBH.
The spectrum evolution of QPEsis also directly related to the evolution of the accretion disk. Future works will require further quantitative analyses of the evolution of mass transfer in combination with specific radiative processes. In addition, continued observations of GSN 096 will help us to understand the interaction of the disk with the star inside it and to test our model.